### v2

# Continual Distillation Learning: An Empirical Study of Knowledge Distillation in Prompt-based Continual Learning

Abstract
Knowledge Distillation (KD) focuses on using a teacher model to improve a student model. Traditionally, KD is studied in an offline fashion, where a training dataset is available before learning. In this work, we introduce the problem of Continual Distillation Learning (CDL) that considers KD in the Continual Learning (CL) setup. A teacher model and a student model need to learn a sequence of tasks, and the knowledge of the teacher model will be distilled to the student to improve the student model in an online fashion. The CDL problem is valuable to study since for prompt-based continual learning methods, using a larger vision transformer (ViT) leads to better performance in continual learning. Distilling the knowledge from a large ViT to a small ViT can improve inference efficiency for prompt-based CL models. To this end, we conducted experiments to study the CDL problem with three prompt-based CL models, i.e., L2P, DualPrompt and CODA-Prompt, where we utilized logit distillation, feature distillation and prompt distillation for knowledge distillation from a teacher model to a student model. Our findings of this study can serve as baselines for future CDL work. [^1] Project website https://irvlutd.github.io/CDL/
1 Introduction
Continual Learning (CL) [30] designs models that can continuously learn new tasks without forgetting previously learned tasks. In class-incremental continual learning, data of new classes arrive sequentially, and the model needs to learn to recognize these classes sequentially during training. In testing, the model will be tested on all the seen classes. Therefore, a good CL model should learn new classes without forgetting.
Traditionally, the class-incremental setting is tackled using a rehearsal buffer [24, 17, 21, 3], which stores training examples from previous classes. During training, data from new tasks and data from the rehearsal buffer are used to update the model weights. With the help of the rehearsal buffer, the model can continually refresh its memory of previous classes. However, due to the size limit of the rehearsal buffer, these models still suffer from the catastrophic forgetting problem [22].
Recently, a new paradigm for CL based on learning for prompt receives more attentions [34, 33, 28, 31]. Instead of using rehearsal buffers, these methods focus on learning prompts for pre-trained large models such as vision transformers for continual learning. For example, L2P [34] learns a prompt pool for a set of tasks using pre-trained ViTs [6], where different tasks learn to use different prompts in the prompt pool. DualPrompt [33] learns two types of prompts, i.e., a global prompt for all tasks and task-specific prompts. By leveraging the pre-trained ViT models, these models achieve better performance compared to traditional rehearsal buffer-based methods.
We noticed that for prompt-based CL methods using ViTs, larger ViTs [6] achieve better performance. For example, ViT large is better than ViT base, and ViT base is better than ViT small, etc. Therefore, it is appealing to use large models in prompt-based CL. However, larger models introduce more computation during inference. To address this problem, we study the problem of Knowledge Distillation (KD) from large to small models in the continual learning setting. We name this problem Continual Distillation Learning (CDL).
Figure 1: Comparison between three different knowledge distillation scenarios. (a) Traditional KD with an offline dataset. (b) KD in rehearsal-based CL models. (c) The CDL problem introduced in this work.
The CDL problem is different from the previous knowledge distillation setup. We illustrate these differences in Fig. 1. First, in the traditional KD setup, an offline dataset is used for knowledge distillation (Fig. 1(a)). A teacher model is first trained using the dataset, and then its knowledge is distilled to a student model using the same dataset. The setup is not the continual learning setup. Second, KD is also used in the rehearsal-based CL methods such as iCaRL [24] (Fig. 1(b)). These methods maintain a rehearsal buffer of previous data. When updating the model parameters for the current task, KD ensures that the prediction on the previous data should be similar to the prediction of the old model. In this case, the old model before updating is the teacher model. Lastly, the CDL problem we study in this work is illustrated in Fig. 1(c). Given a new task in continual learning, a teacher model is first updated. Then a student model is updated based on the data of the new task and the teacher model. Consequently, a better student model can be trained by knowledge distillation from the teacher model.
To study the CDL problem, we have designed a prompt-based continual learning method that consists of a teacher model and a student model. In our method, we jointly learn the teacher model and the student model. Both models are prompt-based CL models, where the backbones are frozen and the prompts are learnable. When training data for new tasks come, the teacher model is first updated as other prompt-based CL models in the literature, and then the student model is updated.
Our method is flexible to be used with different prompt-based CL models and different mechanisms for knowledge distillation. Therefore, we conducted an empirical study of the CDL problem using three prompt-based CL models, i.e., L2P [34], DualPrompt [33] and CODA-Prompt [28]. For each model, we considered different teacher-student configurations. For example, using ViT large as the teacher model and ViT base as the student model. In addition, we studied the effect of three types of knowledge distillation mechanism in the CDL setup, i.e. logit distillation, feature distillation, and prompt distillation. We found that prompt distillation works the best, due to the prompt learning nature of these models.
In summary, our contributions are as follows.
  • We introduce the problem of continual distillation learning (CDL) that studies knowledge distillation of a teacher model to a student model in continual learning.
  • We designed a method for CDL that can be used with different prompt-based CL models and different knowledge distillation mechanism.
  • We conducted extensive experiments and analyses of three prompt-based CL models with three types of KD mechanism on two commonly-used benchmarks for continual learning.
2 Related Work
Continual Learning. The main purpose of continual learning is to build an intelligent system to solve the problem of catastrophic forgetting  [22] in the case of incremental tasks. Different types of continual learning methods have been proposed in the literature [32]. For example, [24, 17] utilize function regularization to help the loss function. There are also weight regularization methods [15, 38, 20, 25, 27] that selectively constrain changes in network parameters and impose penalties on changes in each parameter based on its contribution. Architecture-based approaches [37, 13, 35, 7, 23] mainly focus on constructing a special model for continual learning. Replay-based approaches [24, 17, 12, 21, 1, 3, 2] use the memory buffer to store and replay old task data in learning the current task. Some recent works aim to address the issue of catastrophic forgetting without relying on rehearsal memory, which are referred as rehearsal-free methods [9, 8, 5]. The aforementioned methods do not leverage pre-trained networks. They usually learn the tasks from scratch.
Prompt-based Continual Learning. Inspired by the use of prompts in natural language processing [19], prompt-based continual learning methods employ large Vision Transformer (ViT) models [6]. These methods freeze the pre-trained backbone and shift learning to the prompts of the ViTs. For example, L2P [34] learns a prompt pool of key-prompt pairs, then selects the optimal prompt in the pool for a given input by matching the input with the keys in the pool to select the corresponding prompts. Based on this idea, DualPrompt [33] introduces the concept of using a general prompt and an expert prompt. CODA-Prompt [28] abandons the key-prompt pair selection idea. It uses a weighted sum of the prompt components to obtain the final prompt. In prompt-based methods, the size of the backbone matters. Larger backbone models achieve better performance. This motivates us to study the knowledge distillation problem for prompt-based continual learning methods, where we can distill knowledge from large models to small models in order to speed up inference time.
Knowledge Distillation (KD). The purpose of knowledge distillation is to leverage large models to assist smaller models in fully exploiting their potential. The larger model, termed the teacher, aids in training the smaller model, referred to as the student. According to the distillation location, traditional methods can be divided into logit distillation [11, 39] and feature distillation [18, 26, 4]. With the recent rapid advancements in large models, new research has directed the focus of knowledge distillation towards vision transformers [29, 36]. However, KD for prompt-based continual learning models has not yet been explored, which is the focus of this work.
3 Prerequisites
3.1 L2P
L2P [34] and DualPrompt [33] use a pre-trained ViT model as the backbone and keep the backbone frozen. Both methods use a prompt pool to encode information about tasks: $\mathbf{P}=\{P_{1},P_{2},\ldots,P_{M}\}$, where $P_{i}\in\mathbb{R}^{L_{p}\times D}$ with token length $L_{p}$ and embedding size $D$, and $M$ is the total number of prompts. In order to select prompts for different tasks, each prompt is associated with a learnable key $\{(\mathbf{k}_{1},P_{1}),(\mathbf{k}_{2},P_{2}),\ldots,(\mathbf{k}_{M},P_{M})\}$, where $\mathbf{k}_{i}\in\mathbb{R}^{D_{k}}$ and $\mathbf{K}=\{\mathbf{k}_{i}\}_{i=1}^{M}$ denotes all the keys.
Given an input image $\mathbf{x}$, a query function $q(\cdot)$ is used to encode the input image. The query $\mathbf{x}_{e}=q(\mathbf{x})\in\mathbb{R}^{D_{k}}$ then matches the key $\mathbf{k}_{i}$ with cosine similarity: $\gamma\left(q(x),\mathbf{k}_{i}\right)$, where $\gamma(\cdot,\cdot)$ denotes the cosine similarity function. Then the top $k$ prompts from the prompt pool $\mathbf{P}$ based on the establishment of key-value pairs are selected. These prompts are used in the ViT model to classify the input. In this way, given an input-label pair $(\mathbf{x},y)$, L2P optimizes the trainable prompt keys and values, and the final classification layer. The training loss function is defined as
$\min_{\mathbf{P},\mathbf{K},\phi}\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)+\lambda\sum_{\mathbf{k}_{i}\in\mathbf{K}_{\mathbf{x}}}\gamma(q(\mathbf{x}),\mathbf{k}_{i}),$ (1)
where $g_{\phi}$ is the classifier, $f_{b}$ is the pre-trained ViT backbone, $\mathbf{K}_{\mathbf{x}}$ denotes the selected keys for input $\mathbf{x}$, and $\mathcal{L}$ denotes the softmax cross-entropy loss for classification. The final class token of the ViT is used for the classifier. Therefore, the first part in Eq. (1) is the softmax cross-entropy loss to optimize the learnable prompt set $\mathbf{P}$, key set $\mathbf{K}$ and the classifier parameter $\phi$. The second term minimizes the distances between the selected keys and the corresponding query features $q(\mathbf{x})$.
3.2 DualPrompt
The DualPrompt method [33] supplements insertable prompts based on the L2P method. The prompts are divided into G-Prompt (General) and E-Prompt (Expert). The shared G-Prompt among all tasks and the corresponding E-Prompt are attached to multiple multi-head self-attention (MSA) layers of the pre-trained transformer. The prompts in both the E-Prompt and L2P are identical, primarily serving to distinguish between different tasks. In contrast, the G-Prompt serves as a shared prompt across different tasks, mainly representing the shared information between tasks. In training, DualPrompt optimizes the G-prompt and E-prompt jointly. Please see [33] for details.
3.3 CODA-Prompt
Based on DualPrompt, CODA-Prompt [28] desires an approach that is end-to-end differentiable. Without selecting the top-K key-value pairs in the prompt pool, CODA-Prompt builds the prompt components: $\mathbf{p}=\sum_{m}\alpha_{m}P_{m}$, where $\mathbf{P}=\{P_{1},P_{2},\ldots,P_{M}\}\subseteq\mathbb{R}^{L_{p}\times D\times M}$ is the set of prompt components, $M$ is the length of the set. $\alpha_{m}$ is the weight for the component $P_{m}$, which is computed using the cosine similarity of the Query-Key. Unlike L2P and Dual-Prompt, when calculating the similarity, CODA-Prompt creates a feature-selection attention scheme $\mathbf{A}$ to process the query features: $\alpha=\gamma\left(q(\mathbf{x})\odot\mathbf{A},\mathbf{K}\right)$, where $\mathbf{A}\subseteq\mathbb{R}^{D_{k}\times M}$ are learnable parameters. Finally, CODA-Prompt optimizes the prompt components $\mathbf{P}$, keys $\mathbf{K}$, attention $\mathbf{A}$ and the classifier. Additionally, orthogonality constraints are added to reduce interference between existing and new knowledge. The final total loss function is:
$\min_{\mathbf{P},\mathbf{K},\mathbf{A},\phi}\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)\\ +\lambda\left(\mathcal{L}_{\text{or}}\left(\mathbf{P}\right)+\mathcal{L}_{\text{or}}\left(\mathbf{K}\right)+\mathcal{L}_{\text{or}}\left(\mathbf{A}\right)\right),$ (2)
where $\mathbf{P}$, $\mathbf{K}$, $\mathbf{A}$ refer to the prompt components and corresponding keys/attention vectors during tasks, respectively. The orthogonality penalty loss is $\mathcal{L}_{\text{or}}(B)=\|BB^{T}-I\|_{2}$, where $B$ represents any matrix.
3.4 Prompt Fine-tuning
Based on the characteristics of large models, they possess outstanding feature extraction capabilities. However, due to the large model size, conventional fine-tuning approaches are not particularly effective. Current prompt-based continual learning methods employ a pre-trained Vision Transformer (ViT) as the backbone, with the backbone component frozen, which effectively preserves the internal hidden features of tasks. These methods adopt a prompt fine-tuning approach, shifting the training focus to the prompt pool. This approach only requires tuning a small portion of the model (the prompt embeddings) during the tasks, thereby effectively mitigating the issue of catastrophic forgetting. Based on the manner in which prompts are incorporated, prompt fine-tuning can be categorized into two types: prompt tuning and prefix tuning.
Figure 2: Our teacher-student model for continual distillation learning. The image is processed by the query function to generate a query, which is compared with $\mathbf{K}$ in the prompt pool. Based on different prompt extractor methods, the output prompts are divided into $\mathbf{P_{K}}$ and $\mathbf{P_{V}}$, which are inserted into the MSA layers across multiple blocks. The dashed boxes represent different types of optional distillation methods. Feature distillation uses the internal tokens output by intermediate blocks, and the class-token is used in the classification layer for logit distillation. Prompt Distillation adds an additional KD-Token and KD-Classifier to establish knowledge distillation with the teacher’s classification layer. In this case, the student model’s prediction is influenced by both the KD-Classifier and the Student-Classifier.
Prompt Tuning prepends the prompt to the input tokens. Equivalent to concatenating the prompt parameters to the input query ($\mathbf{h}_{Q}$), key ($\mathbf{h}_{K}$), and values ($\mathbf{h}_{V}$) of the MSA layer, this approach will increase the output length, where the output dimension is $\mathbb{R}^{(L+L_{p})\times D}$:
$f_{\text{Pro-T}}(P,\mathbf{h})=\text{MSA}([P;\mathbf{h}_{Q}],[P;\mathbf{h}_{K}],[P;\mathbf{h}_{V}]).$ (3)
Prefix Tuning divides the prompt $P$ into two parts, $P_{K}$, $P_{V}\in\mathbb{R}^{L_{p}/2\times D}$, which are inserted into the key $\mathbf{h}_{K}$ and value $\mathbf{h}_{V}$, respectively:
$f_{\text{Pre-T}}(P,\mathbf{h})=\text{MSA}(\mathbf{h}_{Q},[P_{K};\mathbf{h}_{K}],[P_{V};\mathbf{h}_{V}]).$ (4)
This method keeps the output sequence length the same as the input. In this paper, since we need to insert prompts across multiple blocks, we employ the prefix method to avoid adding too many additional tokens. Then the prompts retrieved from the prompt pool are divided into two parts, $P_{K}$ and $P_{V}$.
4 Continual Knowledge Distillation
To study the CDL problem, we have designed a teacher-student model based on prompt-based continual learning methods, as illustrated in Fig. 2. Both the teacher model and the student model are some prompt-based continual learner such as L2P, DualPrompt or CODA-Prompt. Typically, the teacher’s backbone utilizes a larger and more accurate pretrained ViT model. The framework supports using different continual learners. We will study three different types of knowledge distillation within the framework: logit distillation, feature distillation and prompt distillation.
4.1 Prompt Extractor
To select a prompt, cosine similarity is used: $\gamma\left(q(\mathbf{x}),\mathbf{k}_{m}\right)$ between the output of a query function $q(\mathbf{x})$ and a prompt key $\mathbf{k}_{m}$. Each key has a corresponding value in the prompt pool, forming key-value pairs. We can consider using different prompt extracting methods for different prompt-based methods. Then the final output prompts are divided into two parts, $P_{K}$ and $P_{V}$. Here, we use Prefix Tuning to insert $P_{K}$ and $P_{V}$ to the ViT backbone.
4.2 Logit Distillation
Logit distillation is one of the most classic forms of knowledge distillation. It aims to have a smaller student model learn the logits output of a larger or more accurate teacher model. In our CDL setup, we conducted experiments using two types of logit knowledge distillation methods: normal knowledge distillation (KD) [11] and Decoupled Knowledge Distillation (DKD) [39].
Normal Knowledge Distillation (KD): This approach was initially proposed by Hinton et al. [11], where distillation transfers the knowledge from a large, complex teacher model to a smaller student model, helping the latter to approximate the teacher model in terms of performance and accuracy. To achieve this, Hinton et al. designed a method where the logit output of the teacher model serves as “soft labels”, guiding the student model’s training. After passing through the softmax layer, the output values provide probabilities for each class.
$p_{i}=\frac{\exp(z_{i}/\tau)}{\sum_{j}\exp(z_{j}/\tau)},$ (5)
where $z_{i}$ is the logit, and $p_{i}$ is the predicted probability for class $i$. The soft target is the class probability distribution of teacher model. In the distillation process, a temperature parameter $\tau$, is introduced to smooth the output distribution of the model, making it easier for the student model to learn the subtle differences between classes.
In the teacher-student model shown in Fig. 2, the class tokens of the teacher model and the student model processed by the pre-trained ViT backbones are connected to the teacher classifier and student classifier to output their logits, respectively. The loss function of the student model for logit distillation is defined as:
$\mathcal{L}_{\text{S}}$ $=(1-\alpha)\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)+\alpha\mathcal{L}_{\text{KD}}+\lambda\mathcal{L}_{\text{pool}},$ (6)
$\mathcal{L}_{\text{KD}}$ $=\tau^{2}\sum_{i}p_{i}^{\mathcal{T}}\log\left(\frac{p_{i}^{\mathcal{T}}}{p_{i}^{\mathcal{S}}}\right),$ (7)
where $\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)$ represents the cross-entropy classification loss used for learning with true label $y$, $g_{\phi}$ is the student classifier, and $f_{b}$ denotes the pre-trained ViT backbone and we only use the final class token into the classifier. $\mathcal{L}_{\text{KD}}$ represents the knowledge distillation loss, i.e., the KL divergence between the teacher’s probability distribution $p_{i}^{\mathcal{T}}$ and the student’s probability distribution $p_{i}^{\mathcal{S}}$. The $\mathcal{L}_{\text{pool}}$ is a loss function specific to the prompt pool. Different prompt-based continual learning methods have their respective prompt pool loss functions. For example, please refer to the loss functions in Equations (1) and (2) for L2P and CODA-Prompt. $\alpha$ and $\lambda$ are hyperparameters used to balance the weights of loss components.
Decoupled Knowledge Distillation (DKD) [39]: Unlike traditional knowledge distillation, which uses a unified KL divergence to measure the difference between the outputs of the student model and the teacher model, DKD separates the logit distillation loss into target class and non-target class components. DKD considers that the target and non-target classes contain different information and should be handled separately during training. By decoupling these components, DKD allows the student model to better capture the confidence on the target class while learning the distribution of the non-target classes. The student model loss function is formulated as follows:
$\mathcal{L}_{\text{S}}$ $=(1-\alpha)\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)$
$+\alpha(\mathcal{L}_{\text{TCKD}}+\mathcal{L}_{\text{NCKD}})+\lambda\mathcal{L}_{\text{pool}},$ (8)
$\mathcal{L}_{\text{TCKD}}$ $=p_{t}^{\mathcal{T}}\log\left(\frac{p_{t}^{\mathcal{T}}}{p_{t}^{S}}\right)+p_{\setminus t}^{\mathcal{T}}\log\left(\frac{p_{t}^{\mathcal{T}}}{p_{\setminus t}^{\mathcal{S}}}\right)$ (9)
$\mathcal{L}_{\text{NCKD}}$ $=p_{\setminus t}^{\mathcal{T}}\sum_{\begin{subarray}{c}i\neq t\end{subarray}}\hat{p}_{i}^{\mathcal{T}}\log\left(\frac{\hat{p}_{i}^{\mathcal{T}}}{\hat{p}_{i}^{\mathcal{S}}}\right)$ (10)
where $\mathcal{L}_{\text{TCKD}}$ is the target class distillation loss, focused on aligning the student’s confidence with the teacher’s for the correct class. $\mathcal{L}_{\text{NCKD}}$ is the non-target class distillation loss, focused on matching the teacher and student model distributions for the incorrect classes. $[p_{t}^{\mathcal{T}},p_{\setminus t}^{\mathcal{T}}]$ represents the binary probabilities of the target class $p_{t}^{T}$ and all the other non-target classes $p_{\setminus t}^{\mathcal{T}}$ in teacher model, which can be calculated by Equation (5). $[p_{t}^{\mathcal{S}},p_{\setminus t}^{\mathcal{S}}]$ represents the binary probabilities in student model. Meanwhile, $\hat{p}_{i}^{\mathcal{T}}$ and $\hat{p}_{i}^{\mathcal{S}}$ are probability distributions among the non-target classes (without considering the $t$th class). Each element is calculated by:
$\hat{p}_{i}=\frac{\exp(z_{i}/\tau)}{\sum\limits_{j\neq t}\exp(z_{j}/\tau)}.$ (11)
4.3 Feature Distillation
Feature distillation focuses on transferring the intermediate representations from a teacher model to a student model. It can leverage richer, layer-wise information within the teacher model to guide the student model’s learning. The student model can benefit from an understanding of feature relationships. In the CDL model in Fig. 2, the backbone ViT consists of multiple blocks, and we use the internal tokens outputted by each block as features. This allows the internal tokens of the student model to learn the information in the internal tokens of the teacher model. It is worth noting that during knowledge distillation, the backbone of the student model remains frozen while processing the internal tokens, which is a characteristic of prompt-based continual learning methods. The frozen part is crucial for mitigating catastrophic forgetting during the continual learning process. In this paper, we study two methods based on the handling of internal tokens: FitNets [26] and Review Knowledge Distillation (ReviewKD) [4].
FitNets [26]: This method is one of the most classic feature distillation methods, aimed at adding a distillation loss to the intermediate layers. It uses the intermediate feature representations of the teacher model as hints to guide the student model’s learning. Here, we select the output of the last block of the teacher model as the hint. Similarly, the student model selects the output of the corresponding block for learning, constructing the feature distillation loss. Finally, the total loss function for the student model is
$\mathcal{L}_{\text{S}}$ $=\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)+\alpha\mathcal{L}_{\text{hint}}+\lambda\mathcal{L}_{\text{pool}},$ (12)
$\mathcal{L}_{\text{hint}}$ $=\left\|f_{b-1}^{\mathcal{T}}(\mathbf{x})-F_{M}(f_{b-1}^{\mathcal{S}}(\mathbf{x}))\right\|^{2},$ (13)
where $\mathcal{L}_{\text{hint}}$ is the feature distillation loss, calculated using the Mean Squared Error (MSE). $\alpha$ and $\lambda$ are used to balance the weights of loss components. $f_{b-1}$ indicates the feature output (internal tokens) after the last block of the ViT model. The student feature is transformed into the same size as the teacher feature with the mapping layer $F_{M}$, which is simply a fully-connected layer in the network.
ReviewKD [4]: It innovates by “reviewing” multiple hidden layers from both the teacher and student models, offering a more comprehensive approach to capturing hierarchical features across the entire model. It reviews the multiple layers utilizing the concept of residual learning. For instance, the feature from $n_{th}$ block of the student is aggregated with the feature from $(n-1)_{th}$ block of the student to mimic the feature from $(n-1)_{th}$ block of the teacher. The total loss function of the student in ReviewKD is
$\mathcal{L}_{\text{S}}=\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)+\alpha\mathcal{L}_{\text{RKD}}+\lambda\mathcal{L}_{\text{pool}},$ (14)
$\mathcal{L}_{\text{RKD}}=\mathcal{D}(\mathbf{F}^{\mathcal{S}}_{n},\mathbf{F}^{\mathcal{T}}_{n})+\sum_{j=n-1}^{1}\mathcal{D}\left(\mathcal{U}(\mathbf{F}_{j}^{\mathcal{S}},\mathbf{F}_{j+1}^{\mathcal{S}}),\mathbf{F}^{\mathcal{T}}_{j}\right),$ (15)
where $\mathcal{L}_{\text{RKD}}$ is the reviewKD loss, and $\mathcal{D}$ is L2 distance between the student features and teacher features. All student features in the equations have passed through the mapping layer $F_{M}$ to make them the same dimension as the teacher features. $\mathbf{F}_{j}^{\mathcal{S}}$ and $\mathbf{F}_{j}^{\mathcal{T}}$ are the features output by the student model and teacher model, respectively, after passing through $j$ blocks. $\mathbf{F}_{j+1}^{\mathcal{S}}$ represents the fused student features at the $(j+1)_{th}$ block. $\mathcal{U}$ is a module used to fuse features, which performs a weighted combination of the two input features. $n$ is the total number of blocks in the ViT backbone. Therefore, in the student model, the fused features obtained at each block are passed to the next higher block to form new feature fusion.
4.4 Prompt Distillation
In the traditional knowledge distillation, the backbone of the student model is trainable, allowing it to learn and approach the teacher model. However, in the prompt-based methods, the backbone is frozen, which limits the update of intermediate features extracted by the student model, making it challenging to effectively learn knowledge from the teacher model’s features. Similarly, the logits are also impacted during training updates.
In continual learning, as tasks increase, the model seeks to retain information from previous tasks. Prompt-based continual learning leverages a pretrained and frozen ViT model to preserve knowledge across tasks. However, in knowledge distillation, the student model aims to be fully trained and adjusted to approach the teacher model. These two objectives are conflicted with each other, which creates mutual interference between the continual learning process and the knowledge distillation process. Subsequently, it affects the performance of the student model. Here, we introduce a new concept Continual Distillation Interference (CDI) to describe this phenomenon. CDI is commonly observed in various continual learning methods that rely on large pretrained models as we will show in our experiments.
To address the continual interference problem of distillation, we explored a prompt distillation method as motivated by DeiT [29]. As shown in Fig. 2, we introduce a learnable prompt called KD-Token in the student model, which operates alongside the class token to assist in classification. The KD-Token in the student model can be considered as learning knowledge from the class token in the teacher model. After passing through the ViT backbone, the KD token is input into a newly added KD classifier. The KD classifier has two roles. First, it is used for knowledge distillation with the teacher model. Second, the output of the KD classifier and the output of the student classifier are averaged to generate the final classification result.
Therefore, the KD-Token is specifically designed to assist in knowledge distillation, while the original prompt tokens $P_{K}$ and $P_{V}$, obtained from the prompt pool, are used to facilitate the continual learning process, each serving its respective role. The purpose of this method is to mitigate the CDI problem in prompt-based continual learning. The training loss function of the student model is the same as the one in Equation (6). However, the KD loss is between the KD classifier and teacher classifier. Here, we highlight the difference by:
$\mathcal{L}_{\text{S}}$ $=(1-\alpha)\mathcal{L}(g_{\phi}^{\mathcal{S}}(f_{b}(\mathbf{x})),y)$
$+\alpha\mathcal{L}_{\text{KD}}(k_{\phi}^{\mathcal{S}}(\mathbf{F}^{\mathcal{S}}_{n}[kd]),g_{\phi}^{\mathcal{T}}(\mathbf{F}^{\mathcal{T}}_{n}[cls]))+\lambda\mathcal{L}_{\text{pool}},$ (16)
where $g_{\phi}^{\mathcal{S}}$ and $k_{\phi}^{\mathcal{S}}$ represent the student classifier and KD classifier in the student model, respectively. $g_{\phi}^{\mathcal{T}}$ is the teacher classifier. $n$ is the total number of blocks in the ViT backbone. $\mathbf{F}^{\mathcal{S}}_{n}[kd]$ represent the KD-Token after the $n_{\text{th}}$ block of the student model. $\mathbf{F}^{\mathcal{T}}_{n}[cls]$ denotes the class token in the teacher model. The knowledge distillation loss $\mathcal{L}_{\text{KD}}$ (Equation (7)) uses KL divergence between the KD classifier and teacher classifier. $\mathcal{L}_{\text{pool}}$ represents the loss function on the prompt pool across different continual learning methods, and $\alpha$ and $\lambda$ are balancing weights.
Methods Split ImageNet-R Split CIFAR-100
Teacher Student Avg. Acc ($\uparrow$) Forgetting ($\downarrow$) Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
$\varnothing$ ViT-Small $67.44\pm 0.46$ $8.52\pm 0.05$ $82.18\pm 0.20$ $6.48\pm 0.48$
ViT-Base ViT-Small KD [11] $69.23\pm 0.05$ $8.57\pm 0.39$ $82.39\pm 0.25$ $7.33\pm 0.21$
ViT-Base ViT-Small DKD [39] $68.92\pm 0.07$ $8.39\pm 0.36$ $82.27\pm 0.20$ $7.81\pm 0.14$
ViT-Base ViT-Small FitNets [26] $69.87\pm 0.04$ $7.38\pm 0.36$ $81.83\pm 0.05$ $8.83\pm 0.48$
ViT-Base ViT-Small ReviewKD [4] $70.19\pm 0.16$ $7.68\pm 0.01$ $82.20\pm 0.41$ $7.54\pm 0.03$
ViT-Base ViT-Small KD-Token [29] $\mathbf{70.94\pm 0.33}$ $\mathbf{6.17\pm 0.07}$ $\mathbf{83.72\pm 0.08}$ $\mathbf{6.40\pm 0.06}$
$\varnothing$ ViT-Base $75.78\pm 0.64$ $5.70\pm 0.66$ $86.16\pm 0.17$ $5.63\pm 0.25$
ViT-Large ViT-Base KD [11] $\mathbf{77.51\pm 0.01}$ $\mathbf{3.44\pm 0.14}$ $87.41\pm 0.05$ $5.45\pm 0.06$
ViT-Large ViT-Base DKD [39] $76.70\pm 0.17$ $4.84\pm 0.12$ $85.42\pm 0.31$ $6.55\pm 0.16$
ViT-Large ViT-Base FitNets [26] $74.55\pm 0.14$ $6.81\pm 0.15$ $85.95\pm 0.25$ $6.56\pm 0.02$
ViT-Large ViT-Base ReviewKD [4] $75.72\pm 0.27$ $4.14\pm 0.05$ $86.21\pm 0.61$ $5.64\pm 0.40$
ViT-Large ViT-Base KD-Token [29] $77.28\pm 0.29$ $4.63\pm 0.16$ $\mathbf{88.10\pm 0.33}$ $\mathbf{5.39\pm 0.19}$
Table 1: The continual knowledge distillation results on the CIFAR-100 dataset and the ImageNet-R dataset with different teacher-student models based on CODA-Prompt [28].
5 Experimental Results
5.1 Experimental Benchmarks and Settings
We utilize the CIFAR-100 [16] and ImageNet-R [10] datasets in a class-incremental continual learning setting for our experiments. Following previous works, we divide the ImageNet-R and CIFAR-100 into 10 tasks, where each task contains 10 classes. We experiment with three prompt-based continual learning models, i.e., L2P [34], DualPrompt [33] and CODAPrompt [28]. In DualPrompt, the G-prompts are set in the first two layers with a prompt length of 5, and the E-prompts are set in layers 3 to 5. CODA-Prompt sets the prompt length to 8, with a total of 100 prompt components inserted into the first 5 layers. This setup is the same for all ViT models. Our experiments tested two teacher-student knowledge distillation pairs. One is distilling from ViT-Large to ViT-Base [6], and the other one is from ViT-Base to ViT-Small [6]. For different knowledge distillation methods, the hyperparameter $\alpha$ for the distillation loss is set to 0.5.
We use two metrics to evaluate the models: accuracy and forgetting rate [21]. The accuracy refers to the average accuracy of all tasks after completing all 10 tasks. The forgetting rate reflects the influence of learning a new task on previously completed tasks. A larger forgetting signifies a more negative impact of the continual learning model. Please see the supplementary material for the details.
5.2 Continual Distillation Results
Methods Split ImageNet-R Split CIFAR-100
Teacher Student Avg. Acc ($\uparrow$) Forgetting ($\downarrow$) Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
$\varnothing$ ViT-Small + L2P [34] $61.95\pm 0.25$ $6.52\pm 0.31$ $77.71\pm 0.49$ $7.12\pm 0.33$
$\varnothing$ ViT-Small + DualPrompt [33] $64.27\pm 0.11$ $5.93\pm 0.03$ $79.85\pm 0.57$ $6.12\pm 0.32$
$\varnothing$ ViT-Small + CODA-Prompt [28] $67.44\pm 0.46$ $8.52\pm 0.05$ $82.18\pm 0.20$ $6.48\pm 0.48$
ViT-Base ViT-Small KD-Token [29] + L2P [34] $65.09\pm 1.05$ $5.37\pm 0.13$ $79.78\pm 0.14$ $7.07\pm 0.52$
ViT-Base ViT-Small KD-Token [29] + DualPrompt [33] $66.81\pm 0.42$ $\mathbf{4.55\pm 0.31}$ $81.32\pm 0.36$ $\mathbf{5.67\pm 0.59}$
ViT-Base ViT-Small KD-Token [29] + CODA-Prompt [28] $\mathbf{70.94\pm 0.33}$ $6.17\pm 0.07$ $\mathbf{83.72\pm 0.08}$ $6.40\pm 0.06$
$\varnothing$ ViT-Base + L2P [34] $71.59\pm 1.20$ $5.65\pm 0.69$ $83.02\pm 0.47$ $6.06\pm 0.47$
$\varnothing$ ViT-Base + DualPrompt [33] $72.44\pm 1.00$ $3.80\pm 1.23$ $84.66\pm 0.87$ $5.91\pm 0.34$
$\varnothing$ ViT-Base + CODA-Prompt [28] $75.78\pm 0.64$ $5.70\pm 0.66$ $86.16\pm 0.17$ $5.63\pm 0.25$
ViT-Large ViT-Base KD-Token [29] + L2P [34] $71.89\pm 0.03$ $\mathbf{3.42\pm 0.36}$ $85.2\pm 0.28$ $\mathbf{4.58\pm 0.55}$
ViT-Large ViT-Base KD-Token [29] + DualPrompt [33] $73.63\pm 0.26$ $3.5\pm 0.25$ $85.32\pm 0.02$ $5.64\pm 0.51$
ViT-Large ViT-Base KD-Token [29] + CODA-Prompt [28] $\mathbf{77.28\pm 0.29}$ $4.63\pm 0.16$ $\mathbf{88.10\pm 0.33}$ $5.39\pm 0.19$
Table 2: Results of prompt distillation using KD token on the ImageNet-R dataset and the CIFAR-100 dataset with different teacher-student models and different continual learning models.
Table 1 summarizes the results of all the aforementioned knowledge distillation methods on the ImageNet-R and CIFAR100 datasets using CODA-Prompt as the continual learner. The complete results for the L2P and DualPrompt methods are provided in the supplementary material. When the teacher model is $\varnothing$, it indicates that there is no knowledge distillation. The student model is trained with the continual learning task only. We hope that the knowledge distillation methods can improve over this baseline by increasing accuracy and reducing forgetting rate.
From the table, we observe that, prompt distillation using KD-Token achieves the best overall performance on the two datasets. The other knowledge distillation methods exhibit certain limitations to varying extents. For example, KD with ViT-Small leads to increased forgetting, and feature distillation methods such as FitNets and ReviewKD even reduce accuracy in some cases. Although KD, DKD, FitNets, and ReviewKD may improve continual learning performance in some situations, their effects are not significant. They cannot simultaneously ensure both accuracy and reduced forgetting. This exemplifies the continual distillation interference problem mentioned in Section  4.4.
In contrast, the prompt distillation method using KD-Token demonstrates a more significant improvement for the student model. Table 2 presents the distillation results of the prompt distillation (KD-Token) method combined with L2P, DualPrompt and CODA-Prompt. We observe that, in all cases, the method with KD-Token distillation outperforms the original model, as reflected in the accuracy and forgetting rate metrics, validating the effectiveness of this approach. In addition, among the three continual learning models, CODA-Prompt achieves the best performance with or without knowledge distillation.
Balance $\mathbf{\alpha}$ 00 $0.1$ $0.3$ $0.5$ $0.7$ $0.9$ $1.0$
Avg. Acc ($\uparrow$) $67.44$ $68.96$ $70.06$ $\mathbf{70.94}$ $69.34$ $68.24$ $68.06$
Forgetting ($\downarrow$) $8.52$ $7.11$ $6.46$ $\mathbf{6.17}$ $6.22$ $6.40$ $6.85$
Table 3: Ablation study of the parameter $\alpha$ based on CODA-Prompt [28] with the Vit-Base to ViT-Small teacher-student model for prompt distillation using a KD-Token.
5.3 Continual Distillation Interference
The prompt distillation method shifts the distillation training process to the KD-Token and KD classifier, thereby enhancing the efficiency of the distillation. Based on our analysis, the freezing of the backbone results in suboptimal training outcomes for logit distillation and feature distillation. However, making the backbone trainable can result in a severe forgetting problem.
To investigate the impact of adding trainable components to the backbone, we conducted a comparative experiment involving the following three models: 1) Normal KD with an additional trainable MLP layer added after the last block of the backbone (Add-MLP); 2) Prompt distillation wiht a KD-Token (Add-KD-Token); and 3) the original model without distillation (No KD). We distill from ViT-Base to ViT-Small using CODA-Prompt. After training on all tasks, we recorded the test accuracy for each task. The results are shown in Fig. 3. It can be observed that KD with an MLP improves the distillation effectiveness for the current task and the most recent task, but results in severe forgetting for earlier tasks. Essentially, the objective of continual learning is to retain information from previous tasks as much as possible (model rigidity), while knowledge distillation aims for the student model to learn from the teacher model as effectively as possible (model adaptability). This introduces a mutual interference between continual learning and distillation, which we refer to as Continual Distillation Interference (CDI). The performance of Add_MLP demonstrates the existence of this problem. The KD-Token method attempts to address this problem by adding a KD-Token, shifting the distillation task to this prompt. We observe that the KD-Token method not only improves the accuracy of the current task through knowledge distillation but also maintains improved performance on previous tasks.
5.4 Ablation Studies
[…]
Figure 3: After completing all tasks, the test accuracy results for each task are recorded. The CDL method uses ViT-Base to ViT-Small, with CODA-Prompt as the baseline.
6 Conclusion and Discussion
We introduce the problem of Continual Distillation Learning (CDL), which considers knowledge distillation in the continual learning setup. We emperically studied three types of knowledge distillation methods in CDL, i.e., logit distillation, feature distillation, and prompt distillation, where three different prompt-based continual learning methods (L2P, DualPrompt, and CODA-Prompt) are utilized for this study. We identified the presence of the Continual Distillation Interference (CDI) phenomenon in the process of CDL. The experimental results demonstrate that the KD-Token method outperforms the other distillation methods, consistently helping the student model improve performance in different scenarios. The primary purpose of this paper is to serve as an inspiration in the hope that more research on CDL will emerge in the future.
Limitations. Knowledge distillation relies on a well-performing teacher model. Prompt-based continual learning methods can obtain teacher models by using large ViT backbones. However, for traditional continual learning methods that rely on CNN backbones (e.g., EWC [15] and iCaRL [24]), it is not straightforward to obtain a well-performing teacher model. This is why we limit our study to prompt-based continual learning. In addition, since each task requires training a larger teacher model, the total training time increases.
Acknowledgments
This work was supported in part by the DARPA Perceptually-enabled Task Guidance (PTG) Program under contract number HR00112220005, the Sony Research Award Program, and the National Science Foundation (NSF) under Grant No. 2346528.
Supplementary Material
{strip}
Methods Split ImageNet-R Split CIFAR-100
Teacher Student Avg. Acc ($\uparrow$) Forgetting ($\downarrow$) Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
$\varnothing$ ViT-Small $61.95\pm 0.25$ $6.52\pm 0.31$ $77.71\pm 0.49$ $7.12\pm 0.33$
ViT-Base ViT-Small KD [11] $63.36\pm 0.04$ $6.74\pm 0.10$ $77.92\pm 0.23$ $8.30\pm 0.29$
ViT-Base ViT-Small DKD [39] $62.91\pm 0.27$ $6.55\pm 0.17$ $78.21\pm 0.12$ $9.13\pm 0.07$
ViT-Base ViT-Small FitNets [26] $64.29\pm 0.09$ $6.37\pm 0.17$ $79.56\pm 0.39$ $\mathbf{5.89\pm 0.36}$
ViT-Base ViT-Small ReviewKD [4] $63.64\pm 0.34$ $6.36\pm 0.58$ $78.50\pm 0.39$ $8.04\pm 0.75$
ViT-Base ViT-Small KD-Token [29] $\mathbf{65.09\pm 1.05}$ $\mathbf{5.37\pm 0.13}$ $\mathbf{79.78\pm 0.14}$ $7.07\pm 0.52$
$\varnothing$ ViT-Base $71.59\pm 1.20$ $5.65\pm 0.69$ $83.02\pm 0.47$ $6.06\pm 0.47$
ViT-Large ViT-Base KD [11] $74.12\pm 0.42$ $4.60\pm 0.55$ $80.96\pm 0.34$ $10.17\pm 0.67$
ViT-Large ViT-Base DKD [39] $\mathbf{74.58\pm 0.01}$ $4.69\pm 0.06$ $83.29\pm 0.24$ $4.99\pm 0.17$
ViT-Large ViT-Base FitNets [26] $70.39\pm 0.23$ $5.84\pm 0.06$ $83.60\pm 0.02$ $5.21\pm 0.71$
ViT-Large ViT-Base ReviewKD [4] $72.17\pm 0.26$ $6.11\pm 0.08$ $83.12\pm 0.65$ $7.97\pm 0.27$
ViT-Large ViT-Base KD-Token [29] $71.89\pm 0.03$ $\mathbf{3.42\pm 0.36}$ $\mathbf{85.20\pm 0.28}$ $\mathbf{4.58\pm 0.55}$
Table 4: The continual knowledge distillation results on the CIFAR-100 dataset and the ImageNet-R dataset with different teacher-student models based on L2P [34].
7 Implementation Details
Our experiments use two metrics to evaluate the models: accuracy and forgetting rate. Accuracy refers to the average accuracy of all tasks after completing all 10 tasks, defined by Equation (17), where $R_{i,j}$ is the test classification accuracy of the model on task $t_{j}$ after observing the last sample from task $t_{i}$. $T$ is the total number of tasks. The forgetting rate (Equation (18)), also known as backward transfer, reflects the influence of learning a new task on previously completed tasks. A higher value signifies a more negative impact of the continual learning model.
$\text{ACC}=\frac{1}{T}\sum_{i=1}^{T}R_{T,i}$ (17)
$\text{Forgetting}=\frac{1}{T-1}\sum_{i=1}^{T-1}R_{T,i}-R_{i,i}$ (18)
For all experiments, we utilized the Adam optimizer [14]. The Split ImageNet-R dataset was trained for 35 epochs, while the Split CIFAR-100 dataset was trained for 20 epochs. The learning rate was set to $l=0.001$. In all loss equations, the balancing parameter was set to $\alpha=0.5$, while the temperature parameter for logit distillation was $\tau=2$. Training was conducted using two NVIDIA A5000 GPUs, each with 24 GB of memory.
8 Continual Distillation Results
Tables Continual Distillation Learning: An Empirical Study of Knowledge Distillation in Prompt-based Continual Learning and 5 present the results of the aforementioned knowledge distillation methods on the ImageNet-R and CIFAR-100 datasets, using L2P [34] and DualPrompt [33] as the continual learners. The results indicate that the KD-Token method consistently optimizes and improves the performance of the original student model across different datasets and various continual learners. Notably, it achieves the highest accuracy in the L2P and DualPrompt models on the CIFAR-10 dataset. Additionally, on the ImageNet-R dataset, the KD-Token method attains the highest accuracy under the base-to-small scheme for both models.
Methods Split ImageNet-R Split CIFAR-100
Teacher Student Avg. Acc ($\uparrow$) Forgetting ($\downarrow$) Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
$\varnothing$ ViT-Small $64.27\pm 0.11$ $5.93\pm 0.03$ $79.85\pm 0.57$ $6.12\pm 0.32$
ViT-Base ViT-Small KD [11] $65.68\pm 0.06$ $7.26\pm 0.29$ $80.16\pm 0.54$ $5.76\pm 0.18$
ViT-Base ViT-Small DKD [39] $65.44\pm 0.05$ $7.27\pm 0.23$ $80.44\pm 0.46$ $6.96\pm 0.44$
ViT-Base ViT-Small FitNets [26] $66.20\pm 0.14$ $5.93\pm 0.17$ $80.70\pm 0.17$ $5.73\pm 0.21$
ViT-Base ViT-Small ReviewKD [4] $65.69\pm 0.91$ $6.56\pm 0.46$ $80.33\pm 0.23$ $5.86\pm 0.53$
ViT-Base ViT-Small KD-Token [29] $\mathbf{66.81\pm 0.42}$ $\mathbf{4.55\pm 0.31}$ $\mathbf{81.32\pm 0.36}$ $\mathbf{5.67\pm 0.59}$
$\varnothing$ ViT-Base $72.44\pm 1.00$ $3.80\pm 1.23$ $84.66\pm 0.87$ $5.91\pm 0.34$
ViT-Large ViT-Base KD [11] $73.90\pm 0.14$ $\mathbf{3.31\pm 0.04}$ $83.4\pm 0.74$ $8.23\pm 1.13$
ViT-Large ViT-Base DKD [39] $\mathbf{75.24\pm 0.33}$ $4.15\pm 0.23$ $84.93\pm 0.16$ $\mathbf{4.95\pm 0.12}$
ViT-Large ViT-Base FitNets [26] $71.23\pm 0.04$ $5.71\pm 0.55$ $83.12\pm 0.86$ $8.33\pm 1.36$
ViT-Large ViT-Base ReviewKD [4] $72.19\pm 0.01$ $5.72\pm 0.20$ $84.11\pm 0.92$ $5.19\pm 0.33$
ViT-Large ViT-Base KD-Token [29] $73.63\pm 0.26$ $3.50\pm 0.25$ $\mathbf{85.32\pm 0.02}$ $5.64\pm 0.51$
Table 5: The continual knowledge distillation results on the CIFAR-100 dataset and the ImageNet-R dataset with different teacher-student models based on DualPrompt [33].