### v3

# Continual Distillation Learning: Knowledge Distillation in Prompt-based Continual Learning

Abstract
We introduce the problem of continual distillation learning (CDL) in order to use knowledge distillation (KD) to improve prompt-based continual learning (CL) models. The CDL problem is valuable to study since the use of a larger vision transformer (ViT) leads to better performance in prompt-based continual learning. The distillation of knowledge from a large ViT to a small ViT can improve the inference efficiency for prompt-based CL models. We empirically found that existing KD methods such as logit distillation and feature distillation cannot effectively improve the student model in the CDL setup. To this end, we introduce a novel method named Knowledge Distillation based on Prompts (KDP), in which globally accessible prompts specifically designed for knowledge distillation are inserted into the frozen ViT backbone of the student model. We demonstrate that our KDP method effectively enhances the distillation performance in comparison to existing KD methods in the CDL setup. [^1] Project website https://irvlutd.github.io/CDL/
1 Introduction
Continual Learning (CL) [32] designs models that can continuously learn new tasks without forgetting previously learned tasks. For example, in class-incremental continual learning [26], the data of new classes arrive sequentially, and the model needs to learn to recognize these classes sequentially during training. In testing, the model will be tested on all the seen classes. Therefore, a good CL model should learn new classes without forgetting.
With recent advances in vision models, vision transformers (ViTs) [6] have demonstrated their advantages over convolutional neural networks (CNNs). Similarly, continual learning research has evolved from traditional CNN-based methods (e.g., using ResNet [11]) to the latest prompt-based CL methods that leverage ViTs as the backbone, which now represent the state-of-the-art CL models. Unlike traditional CL models where the backbone is trained, prompt-based CL methods such as L2P [35], DualPrompt [34] and CODA-Prompt [30], utilize a pre-trained ViT as the backbone while keeping it frozen during training. Instead, they optimize a prompt pool designed for continual learning. Using pre-trained ViT models and shifting learning to prompts, these methods achieve better performance compared to CNN-based continual learning approaches such as iCaRL [26] and LWF [19].
Figure 1: Comparison between two knowledge distillation scenarios. (a) Traditional KD with an offline dataset. (b) The CDL problem introduced in this work.
We noticed that for prompt-based CL methods using ViTs, larger ViTs [6] achieve better performance. For example, ViT large is better than ViT base, and ViT base is better than ViT small, etc. Therefore, it is appealing to use large models in prompt-based CL. However, larger models introduce more computation during inference. To address this limitation, we propose a new research problem: Knowledge Distillation (KD) from large to small models in prompt-based continual learning. We name this problem Continual Distillation Learning (CDL), which aims to improve the performance of small models by using large models as teachers in continual learning.
We particularly limit the scope of this work to prompt-based CL models, since we have experimentally verified that large models may not result in better performance for the traditional CNN-based CL models such as the iCaRL [26] model or the LWF [19] model (see Sec. 7 in supplementary material). These CNN-based models update the backbones during training. Deeper and larger CNNs tend to overfit new tasks, failing to fundamentally solve the catastrophic forgetting problem in continual learning.
The CDL problem is different from the previous knowledge distillation setup. We illustrate the differences in Fig. 1. In the traditional KD setup [13], an offline dataset is used for knowledge distillation (Fig. 1(a)). A teacher model is first trained using the dataset, and then its knowledge is distilled to a student model using the same dataset. The setup is not the continual learning setup. The CDL problem that we study in this work is illustrated in Fig. 1(b). Given a new task in continual learning, a teacher model is first updated. Then a student model is updated based on the data of the new task and the teacher model. Consequently, a better student model can be trained by distilling knowledge from the teacher model. In general, this is a task-incremental learning process, where both the teacher and student models can only access the data for the current task at each step.
To explore the new CDL problem, we first conducted an empirical study by applying existing knowledge distillation methods, i.e., logit distillation [13, 40], feature distillation [28, 4] and distillation token [31], to three prompt-based continual learning models, i.e., L2P [35], DualPrompt [34], and CODA-Prompt [30]. For each model, we consider different teacher-student configurations. For example, we can use ViT-Large as the teacher model and ViT-Base as the student model. Since the ViT backbone is frozen in prompt-based CL, knowledge transfer from the teacher to the student can only be based on the prompt pool, where the selection of prompts relies on a key-query mechanism. However, for knowledge distillation, this mechanism is not effective. Since each task independently reselects prompt components from the prompt pool, the knowledge embedded within them from the teacher model cannot be transferred to the next task. Our experimental results also show that the performance improvements achieved by these knowledge distillation methods are not significant. Previous KD methods are not effective in the CDL setting.
To address this limitation, we propose a novel method named Knowledge Distillation based on Prompts (KDP). KDP inserts globally accessible prompts, specifically designed for knowledge distillation, into the frozen ViT backbone of the student model. We name them KD prompts. These prompts are independent of the key-query mechanism of the prompt pool. They are not limited to specific tasks and can facilitate cross-task distillation. They serve as auxiliary information to guide the learning process of the student model. Experimental results demonstrate that this novel approach effectively addresses the CDL problem and enhances distillation performance compared to previous KD methods. Furthermore, we conducted an ablation study on the total number of KD prompts inserted into the backbone. The results show that the introduction of KD prompts into each ViT block significantly improves the alignment of the feature structures between layers of the teacher model and the student model, leading to better performance of the student model.
In summary, our contributions are as follows.
  • We introduce the new problem of continual distillation learning (CDL) that studies knowledge distillation of a teacher model to a student model in continual learning.
  • We conducted an empirical study by applying different knowledge distillation methods to three prompt-based CL approaches and identify the limitations of these distillation methods in the CDL setting.
  • To address the unique challenges of the CDL problem, we propose a novel distillation method, KDP, which outperforms previous KD approaches.
2 Related Work
2.1 Continual Learning
The main purpose of continual learning is to build an intelligent system to solve the problem of catastrophic forgetting [24] in the case of incremental tasks. Different types of continual learning methods have been proposed in the literature [33]. For example, [26, 19] utilize function regularization to help the loss function. There are also weight regularization methods [17, 39, 22, 27, 29] that selectively constrain changes in network parameters and impose penalties on changes in each parameter based on its contribution. Architecture-based approaches [38, 15, 36, 7, 25] mainly focus on constructing a special model for continual learning. Replay-based approaches [26, 19, 14, 23, 1, 3, 2] use a memory buffer to store and replay old task data in learning the current task. Some recent works aim to address the issue of catastrophic forgetting without relying on rehearsal memory, which are referred as rehearsal-free methods [10, 8, 5]. The above methods are primarily based on CNN backbones, and a larger backbone model may not achieve better performance.
2.2 Prompt-based Continual Learning
Inspired by the use of prompts in natural language processing [21], prompt-based continual learning methods employ large vision transformer (ViT) models [6]. These methods freeze the pre-trained backbone and shift learning to prompts of the ViTs. For example, L2P [35] learns a prompt pool of key-prompt pairs, then selects the optimal prompt in the pool for a given input by matching the input with the keys in the pool. Based on this idea, DualPrompt [34] introduces the concept of using a general prompt and an expert prompt. CODA-Prompt [30] abandons the key-prompt pair selection idea. It uses a weighted sum of the prompt components to obtain the final prompt. CPrompt [9] introduces two main components during training: Classifier Consistency Learning (CCL) and Prompt Consistency Learning (PCL). In prompt-based methods, the size of the backbone matters. Larger backbone models achieve better performance. This motivates us to study the knowledge distillation problem for prompt-based continual learning methods, where we can distill knowledge from large models to small models in order to improve small models with fast inference time.
2.3 Knowledge Distillation
The purpose of knowledge distillation is to leverage larger models to assist smaller models to fully exploit their potential. The larger model, termed the teacher, aids in training the smaller model, termed the student. Traditional KD methods can be classified into logit distillation [13, 40] and feature distillation [20, 28, 4]. With recent rapid advancements in large models, new research has directed the focus of knowledge distillation towards vision transformers [31, 37]. However, KD for prompt-based continual learning models has not yet been explored, which is the focus of this work.
3 Prerequisites
3.1 Continual Learning Setting
In continual learning, a model is required to learn a sequence of tasks, where the data from these tasks arrive on time. We denote a sequence of tasks as $\mathcal{D}=\{\mathcal{D}_{1},\ldots,\mathcal{D}_{T}\}$, where $T$ is the number of tasks. The $t$th task $\mathcal{D}_{t}=\{(\mathbf{x}_{i}^{t},y_{i}^{t})\}_{i=1}^{n_{t}}$ consists of pairs of input sample $\mathbf{x}_{i}^{t}\in\mathcal{X}$ and its label $y_{i}^{t}\in\mathcal{Y}$, where $n^{t}$ is the number of samples for the $t$th task. In this work, we consider class-incremental learning, where each task consists of a fixed number of non-overlapping classes. In training, a model learns these tasks one by one. Data from the previous tasks are not available anymore when training future tasks. In testing, the model is evaluated by testing samples from all classes.
Figure 2: Our teacher-student model for continual distillation learning. The green and red dashed boxes represent different types of optional distillation methods. The blue dashed box represents our KDP method, where the components inside are the trainable modules to be added.
3.2 Prompt-based Methods
Prompt-based continual learning methods use pre-trained ViT backbones and freeze the backbones during training. Learning is shifted to trainable prompts.
L2P: L2P [35] uses a prompt pool to encode information about tasks: $\mathcal{P}=\{\mathbf{P}_{1},\mathbf{P}_{2},\ldots,\mathbf{P}_{M}\}$, where $\mathbf{P}_{i}\in\mathbb{R}^{L_{p}\times D}$ with token length $L_{p}$ and embedding size $D$, and $M$ is the total number of prompt components. To select prompts for different tasks, each prompt component is associated with a learnable key $\{(\mathbf{k}_{1},\mathbf{P}_{1}),(\mathbf{k}_{2},\mathbf{P}_{2}),\ldots,(\mathbf{k}_{M},\mathbf{P}_{M})\}$, where $\mathbf{k}_{i}\in\mathbb{R}^{D_{k}}$ with embedding size $D_{k}$ and $\mathcal{K}=\{\mathbf{k}_{i}\}_{i=1}^{M}$ denotes all the keys. Given an input image $\mathbf{x}\in\mathcal{X}$, a query function $q(\cdot)$ is used to encode the input image. The query $\mathbf{x}_{e}=q(\mathbf{x})\in\mathbb{R}^{D_{k}}$ then matches the key $\mathbf{k}_{i}$ with cosine similarity: $\gamma\left(q(\mathbf{x}),\mathbf{k}_{i}\right)$, where $\gamma(\cdot,\cdot)$ denotes the cosine similarity function. Then the top $K$ prompt components from the prompt pool $\mathcal{P}$ based on the establishment of key-value pairs are selected:
$\mathbf{P}_{\mathbf{x}}=\text{TopK}(\mathbf{P}_{1},\mathbf{P}_{2},\ldots,\mathbf{P}_{M}),$ (1)
where TopK indicates the selection function based on cosine similarity. $\mathbf{P}_{\mathbf{x}}\in\mathbb{R}^{K\times L_{p}\times D}$ is the final set of prompt components extracted from the prompt pool to assist continual learning. The training loss function for a pair $(\mathbf{x},y)$ is defined as:
$\min_{\mathcal{P},\mathcal{K},\phi}\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)+\lambda\sum_{\mathbf{k}_{i}\in\mathbf{K}_{\mathbf{x}}}\gamma(q(\mathbf{x}),\mathbf{k}_{i}),$ (2)
where $g_{\phi}$ is the classifier with parameter $\phi$. $f_{b}$ is the pre-trained ViT backbone, which includes the selected prompt components $\mathbf{P}_{\mathbf{x}}$ for continual learning. $\mathbf{K}_{\mathbf{x}}$ denotes the selected keys for input $\mathbf{x}$, and $\mathcal{L}$ denotes the softmax cross-entropy loss for classification. The final class token of the ViT is used for the classifier. Therefore, the first term in Eq. (2) is the softmax cross-entropy loss to optimize the learnable prompt set $\mathcal{P}$ and the classifier parameter $\phi$. The second term learns the key set $\mathcal{K}$ by minimizing the distances between the selected keys and the corresponding query features $q(\mathbf{x})$.
DualPrompt: The DualPrompt method [34] supplements insertable prompts based on the L2P [35] method. The prompts are divided into G-Prompt (General) and E-Prompt (Expert). The shared G-Prompt among all tasks and the corresponding E-Prompt are attached to multiple multi-head self-attention (MSA) layers of the pre-trained transformer. The prompts in both the E-Prompt and L2P are identical, primarily serving to distinguish between different tasks. In contrast, the G-Prompt mainly represents the shared information between tasks. In training, DualPrompt optimizes the G-prompt and E-prompt jointly.
CODA-Prompt: Instead of selecting the top $K$ key-value pairs in the prompt pool, CODA-Prompt [30] uses a weighted summation over the prompt components to compute the learnable prompt parameter for an input $\mathbf{x}$:
$\mathbf{P}_{\mathbf{x}}=\sum_{i=1}^{M}\alpha_{i}\mathbf{P}_{i},$ (3)
where $\alpha_{i}$ is the weight for the prompt component $\mathbf{P}_{i}$, which is computed using the cosine similarity of the query and key. Unlike L2P and DualPrompt, when calculating the similarity, CODA-Prompt creates a feature-selection attention scheme $\mathcal{A}=\{\mathbf{A}_{1},\ldots,\mathbf{A}_{M}\}$ for $M$ prompt components to process the query features, where $\mathbf{A}_{i}\in\mathbb{R}^{D_{k}},i=1,\ldots,M$. The weight is computed as $\alpha_{i}=\gamma\left(q(\mathbf{x})\odot\mathbf{A}_{i},\mathbf{k}_{i}\right)$, for $i=1,\ldots,M$. Finally, CODA-Prompt optimizes the prompt components $\mathcal{P}$, keys $\mathcal{K}$, attention $\mathcal{A}$ and the classifier. Additionally, orthogonality constraints are added to reduce interference between existing and new knowledge. The final loss function is:
$\min_{\mathcal{P},\mathcal{K},\mathcal{A},\phi}\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)\\ +\lambda\left(\mathcal{L}_{\text{or}}\left(\mathcal{P}\right)+\mathcal{L}_{\text{or}}\left(\mathcal{K}\right)+\mathcal{L}_{\text{or}}\left(\mathcal{A}\right)\right),$ (4)
where $\mathcal{P}$, $\mathcal{K}$, $\mathcal{A}$ refer to the prompt set and corresponding keys and attention vectors during tasks, respectively. The orthogonality penalty loss is defined as $\mathcal{L}_{\text{or}}(B)=\|BB^{T}-I\|_{2}$, where $B$ represents any matrix and $I$ is the identity matrix.
3.3 Prompting with Prefix-Tuning
After computing the prompt $\mathbf{P}\in\mathbb{R}^{L_{p}\times D}$ for an input sample, we need to insert this prompt into the ViT backbone for prompting. We use the prefix-tuning method. It divides the prompt $\mathbf{P}$ into two parts, $\mathbf{P}_{K}$, $\mathbf{P}_{V}\in\mathbb{R}^{L_{p}/2\times D}$, which are inserted into the key $\mathbf{h}_{K}$ and value $\mathbf{h}_{V}$ of the multi-head self-attention (MSA) layers, respectively:
$f_{\text{Pre-T}}(\mathbf{P},\mathbf{h})=\text{MSA}(\mathbf{h}_{Q},[\mathbf{P}_{K};\mathbf{h}_{K}],[\mathbf{P}_{V};\mathbf{h}_{V}]).$ (5)
This method keeps the output sequence length the same as the input. In the CDL setup, since we need to insert prompts across multiple blocks, we employ the prefix-tuning method to avoid adding too many additional tokens.
4 Continual Distillation Learning
4.1 Teacher-Student Model
To study the CDL problem, we propose a teacher-student model setup based on prompt-based continual learning methods, as illustrated in Fig. 2. Both the teacher model and the student model are prompt-based continual learners, such as L2P [35], DualPrompt [34], or CODA-Prompt [30]. Typically, the teacher’s backbone utilizes a larger and more accurate pre-trained ViT model. The framework supports using different continual learners.
As illustrated in Fig. 2, an input image $\mathbf{x}$ is first processed by the query function to generate a query, which is compared with keys $\mathcal{K}$ in the prompt pool. Different prompt-based continual learners select the prompts $\mathbf{P}^{\text{cl}}$ for continual learning through the Prompt Extractor, following Eq. (1) or Eq. (3). The prompts are divided into $\mathbf{P}^{\text{cl}}_{K}$ and $\mathbf{P}^{\text{cl}}_{V}$, which are inserted into the MSA layers across multiple blocks with prefix-tuning as in Eq. (5). This process is applied both the teacher and the student. Within this framework, we study and build different types of knowledge distillation in continual learning, including logit distillation and feature distillation.
Logit distillation is one of the most classic forms of knowledge distillation. It aims to have a smaller student model learn the logits output of a larger or more accurate teacher model. In our CDL setup, we built two CDL methods based on the processing of logits: normal knowledge distillation (KD) [13] and Decoupled Knowledge Distillation (DKD) [40].
Feature distillation focuses on transferring the intermediate representations from a teacher model to a student model. In Fig. 2, the backbone ViT consists of multiple blocks, and we use the internal tokens outputted by each block as features. This allows the internal tokens of the student model to learn the information in the internal tokens of the teacher model. We built two CDL methods based on the handling of internal tokens: FitNets [28] and Review Knowledge Distillation (ReviewKD) [4].
Fig. 2 illustrates the logit distillation and the feature distillation in our continual distillation learning scenario.
Training of the teacher-student model. First, the images of the current task are trained on the teacher model. The teacher model is then used to help train the student model. During the student model training process, the teacher model remains frozen and is only used in the inference phase to provide soft labels (logit distillation) or internal tokens (feature distillation) to the student.
4.2 Limitations of Existing KD Methods in CDL
Although we can apply previous knowledge distillation methods in the CDL setting, there are inherent limitations when applying these methods to prompt-based CL. The primary issue arises from the query-key mechanism used in the prompt selection process within prompt-based CL methods.
In the prompt pool, different prompt components can be interpreted as distinct feature representations. Through the query-key mechanism, images from different tasks can select different types of prompt feature combinations, where each image has its own preferred prompt components. The final prompts inserted into the backbone are either these selective prompt components (L2P, DualPrompt) or their weighted combinations (CODA-Prompt).
In prompt-based CL, a new task may select prompt components different from prior tasks, which is an effective strategy to prevent forgetting in continual learning. Because each task can use its own prompts from the prompt pool. However, for knowledge distillation, this mechanism is not effective. For example, in task $A$, a subset of prompt components $\mathcal{P}_{A}$ of the student model is used for learning and knowledge distillation. We hypothesize that these prompt components of the student $\mathcal{P}_{A}$ inherit certain information for the teacher. When task $B$ comes, another subset of prompt components $\mathcal{P}_{B}$ of the student model is selected for learning and knowledge distillation. In the extreme case, if $\mathcal{P}_{A}\bigcap\mathcal{P}_{B}=\emptyset$, the prompts in $\mathcal{P}_{B}$ has no information about the teacher model. Consequently, the student needs to learn from scratch about the teacher model for task $B$.
In knowledge distillation, the goal of the student model is to learn the overall structural knowledge of the teacher model. Distillation should be a global process rather than being confined independently to each task.
4.3 Knowledge Distillation based on Prompts
To overcome the limitations of previous KD methods in CDL, we introduce a novel knowledge distillation method named Knowledge Distillation based on Prompts (KDP), which is specifically designed for the CDL problem. We introduce a new type of prompts named KD prompts that are globally shared between tasks and are specifically designed for knowledge distillation as illustrated in Fig. 2.
Since these components are shared across all tasks, they do not suffer from the limitation mentioned before. Moreover, KD prompts do not rely on the query-key mechanism and do not require selection among prompt components. When a new task comes, knowledge is continuously transferred from the teacher to the student using the KD prompts. Furthermore, introducing additional KD prompts increases the number of learnable prompt components for the student model, preventing all distillation-related updates from being concentrated within the original prompt pool. Consequently, the prompt components of the student model are divided into two categories: the original prompts for continual learning (CL prompts) and the distillation prompts for learning from the teacher model (KD prompts).
Similar to the CL prompts, KD prompts are inserted into the ViT backbone using the prefix-tuning method as in Eq. (5). Specifically, a KD prompt $\mathbf{P}^{\text{kd}}\in\mathbb{R}^{L_{p}\times D}$ is divided into two parts: $\mathbf{P}^{\text{kd}}_{K}$, $\mathbf{P}^{\text{kd}}_{V}\in\mathbb{R}^{L_{p}/2\times D}$, which are inserted into the key $\mathbf{h}_{K}$ and value $\mathbf{h}_{V}$ at the end of each MSA layer. The prefix-tuning for our KDP method is defined as
$f_{\text{Pre-T}}(\mathbf{P}^{\text{cl}},\mathbf{h},\mathbf{P}^{\text{kd}})$
$=\text{MSA}(\mathbf{h}_{Q},[\mathbf{P}^{\text{cl}}_{K};\mathbf{h}_{K},\mathbf{P}^{\text{kd}}_{K}],[\mathbf{P}^{\text{cl}}_{V};\mathbf{h}_{V},\mathbf{P}^{\text{kd}}_{V}]).$ (6)
For processing distilled knowledge, we adopt the distillation through attention method in DeiT [31]. In the student model, a distillation token (the KD token in Fig. 2) is inserted at the end of the first layer of the ViT backbone. After being propagated through the ViT backbone, the final output of the distillation embeddings of the KD token are connected to a separate classifier named the KD classifier.
During training, the KD classifier is used to compute the loss function with the teacher classifier using the normal logit distillation loss function. That is, the logits output of the teacher model serves as the “soft labels” to the student. The target is the softened probability distribution of the teacher model, controlled by the temperature factor $\tau$. Therefore, the loss function for the distillation process is:
$\mathcal{L}_{\text{KD}}=\tau^{2}\sum_{i}p_{i}^{\mathcal{T}}\log\left(\frac{p_{i}^{\mathcal{T}}}{p_{i}^{\mathcal{S}}}\right),$ (7)
where $\mathcal{L}_{\text{KD}}$ is the KL divergence between between the teacher’s probability distribution $p_{i}^{\mathcal{T}}$ and the student’s probability distribution $p_{i}^{\mathcal{S}}$. $\tau$ is the temperature parameter.
The final training loss function of the student model is:
$\mathcal{L}_{\text{S}}$ $=\underbrace{(1-\alpha)\mathcal{L}(g_{\phi}^{\mathcal{S}}(f_{b}^{\mathcal{S}}(\mathbf{x};\mathbf{P}^{\text{kd}}_{1}:\mathbf{P}^{\text{kd}}_{n})),y)}_{\text{student classification loss}}$
$+\underbrace{\alpha\mathcal{L}_{\text{KD}}(k_{\phi}^{\mathcal{S}}(f^{\mathcal{S}}_{b}(\mathbf{x};\mathbf{P}^{\text{kd}}_{1}:\mathbf{P}^{\text{kd}}_{n})),g_{\phi}^{\mathcal{T}}(f^{\mathcal{T}}_{b}(\mathbf{x})))}_{\text{student knowledge distillation loss}}$
$+\underbrace{\lambda\mathcal{L}_{\text{pool}}}_{\text{student prompt pool loss}},$ (8)
where $g_{\phi}^{\mathcal{S}}$ and $k_{\phi}^{\mathcal{S}}$ represent the student classifier and KD classifier in the student model, respectively. $g_{\phi}^{\mathcal{T}}$ is the teacher classifier. $f_{b}^{\mathcal{S}}$ and $f_{b}^{\mathcal{T}}$ represent the ViT backbones of the student and teacher models, respectively, which already include the original CL prompts. $\mathbf{P}^{\text{kd}}_{n}$ denotes the KD prompt inserted into the $n$-th block of the ViT backbone. In total, KD prompts are inserted from the first to the $n$-th block. $\mathcal{L}_{\text{pool}}$ is the loss function involved in extracting CL prompts from the prompt pool. Different prompt-based CL methods have different $\mathcal{L}_{\text{pool}}$, as shown in the latter parts of Eq. (2) and Eq. (4). $\alpha$ and $\lambda$ are balancing weights. During testing, the KD classifier and the student classifier are both used for prediction.
Methods Split ImageNet-R Split CIFAR-100
Teacher Student Avg. Acc ($\uparrow$) Forgetting ($\downarrow$) Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
$\varnothing$ ViT-Small $67.44\pm 0.46$ $8.52\pm 0.05$ $82.18\pm 0.20$ $6.48\pm 0.48$
ViT-Base ViT-Small KD [13] $69.91\pm 0.62$ $7.64\pm 0.71$ $83.03\pm 0.39$ $7.24\pm 0.30$
ViT-Base ViT-Small DKD [40] $68.92\pm 0.07$ $8.39\pm 0.36$ $82.27\pm 0.20$ $7.81\pm 0.14$
ViT-Base ViT-Small FitNets [28] $69.87\pm 0.04$ $7.38\pm 0.36$ $81.83\pm 0.05$ $8.83\pm 0.48$
ViT-Base ViT-Small ReviewKD [4] $70.19\pm 0.16$ $7.68\pm 0.01$ $82.20\pm 0.41$ $7.54\pm 0.03$
ViT-Base ViT-Small DeiT [31] $70.74\pm 0.20$ $6.66\pm 0.28$ $83.79\pm 0.15$ $6.58\pm 0.06$
ViT-Base ViT-Small KDP (ours) $\mathbf{71.92\pm 0.50}$ $\mathbf{5.61\pm 0.34}$ $\mathbf{84.31\pm 0.01}$ $\mathbf{5.63\pm 0.02}$
$\varnothing$ ViT-Base $76.42\pm 0.17$ $4.31\pm 0.18$ $86.16\pm 0.17$ $5.63\pm 0.25$
ViT-Large ViT-Base KD [13] $76.99\pm 0.02$ $3.81\pm 0.06$ $86.27\pm 0.05$ $5.45\pm 0.06$
ViT-Large ViT-Base DKD [40] $76.70\pm 0.17$ $4.84\pm 0.12$ $85.42\pm 0.31$ $6.55\pm 0.16$
ViT-Large ViT-Base FitNets [28] $74.55\pm 0.14$ $6.81\pm 0.15$ $85.95\pm 0.25$ $6.56\pm 0.02$
ViT-Large ViT-Base ReviewKD [4] $75.72\pm 0.27$ $4.14\pm 0.05$ $86.21\pm 0.61$ $5.64\pm 0.40$
ViT-Large ViT-Base DeiT [31] $77.83\pm 0.55$ $4.51\pm 0.03$ $86.78\pm 0.15$ $5.43\pm 0.25$
ViT-Large ViT-Base KDP (ours) $\mathbf{78.62\pm 0.57}$ $\mathbf{3.46\pm 0.53}$ $\mathbf{87.13\pm 0.09}$ $\mathbf{5.30\pm 0.06}$
Table 1: The continual knowledge distillation results on the CIFAR-100 dataset and the ImageNet-R dataset with different teacher-student models based on CODA-Prompt [30].
5 Experiments
5.1 Datasets
We utilize the CIFAR-100 [18] and ImageNet-R [12] datasets in a class-incremental continual learning setting for our experiments. Following previous works, we divide the ImageNet-R and CIFAR-100 datasets into 10 tasks, where each task contains 10 classes.
5.2 Implementation Details
We experiment with three prompt-based continual learning models, i.e., L2P [35], DualPrompt [34] and CODAPrompt [30]. In L2P, the prompt pool consists of a total of 30 prompt components, and a CL prompt with a length of 20 is inserted only in the first layer of the ViT backbone. In DualPrompt, the prompt pool contains 10 prompt components, each CL prompt has a length of 20, with G-prompts and E-prompts placed in the first two blocks and the third to fifth blocks, respectively. In CODA-Prompt, the prompt pool consists of 100 prompt components. Each CL prompt has a length of 8, and CL prompts are inserted from the first to the fifth layer of the ViT. In our KDP method, unless otherwise specified, all KD prompts have a length of 6 and are inserted from the first block layer to the twelfth layer. Our KDP method by default adopts the DeiT method [31] (including both KD token and the KD classifier) to obtain predictions. This is an optional component that can be added or removed. We evaluate the results of removing the DeiT structure from KDP in the ablation study Sec. 5.5. The parameter $\alpha$ used in Eq. (8) to balance distillation and continual learning is set to 0.5. $\lambda$ for $\mathcal{L}_{\text{pool}}$ is set to 1.
In the experiments, all models are trained only on the data of the current task, and the distillation is based on rehearsal-free models. Our experiments mainly tested two teacher-student knowledge distillation pairs. One is distilling from ViT-Large to ViT-Base [6], and the other one is from ViT-Base to ViT-Small [6].
For all experiments, we utilized the Adam optimizer [16]. The Split ImageNet-R dataset was trained for 35 epochs, while the Split CIFAR-100 dataset was trained for 20 epochs. The learning rate was set to $l=0.001$. In all loss equations, the balancing parameter was set to $\alpha=0.5$, while the temperature parameter for logit distillation was $\tau=2$. Training was conducted using two NVIDIA A5000 GPUs, each with 24 GB of memory.
5.3 Evaluation Metrics
Our experiments use two metrics to evaluate the models: accuracy and forgetting rate [23]. Accuracy refers to the average accuracy of all tasks after completing all 10 tasks, defined by Eq. (9), where $R_{i,j}$ is the test classification accuracy of the model on task $t_{j}$ after observing the last sample from task $t_{i}$. $T$ is the total number of tasks. The forgetting rate (Eq. (10)), also known as backward transfer, reflects the influence of learning a new task on previously completed tasks. A higher value signifies a more negative impact of the continual learning model.
$\text{ACC}=\frac{1}{T}\sum_{i=1}^{T}R_{T,i},$ (9)
$\text{Forgetting}=\frac{1}{T-1}\sum_{i=1}^{T-1}R_{T,i}-R_{i,i}.$ (10)
5.4 Continual Distillation Results
Methods Split ImageNet-R Split CIFAR-100
Teacher Student Avg. Acc ($\uparrow$) Forgetting ($\downarrow$) Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
$\varnothing$ ViT-Small + L2P [35] $63.82\pm 0.25$ $6.52\pm 0.31$ $77.71\pm 0.49$ $7.12\pm 0.33$
$\varnothing$ ViT-Small + DualPrompt [34] $65.51\pm 0.11$ $5.93\pm 0.03$ $79.85\pm 0.57$ $6.12\pm 0.32$
$\varnothing$ ViT-Small + CODA-Prompt [30] $67.44\pm 0.46$ $8.52\pm 0.05$ $82.18\pm 0.20$ $6.48\pm 0.48$
ViT-Base ViT-Small KDP (ours) + L2P [35] $68.18\pm 0.03$ $\mathbf{2.08\pm 0.28}$ $81.79\pm 0.66$ $4.31\pm 0.27$
ViT-Base ViT-Small KDP (ours) + DualPrompt [34] $68.77\pm 0.16$ $3.13\pm 0.25$ $81.78\pm 0.17$ $\mathbf{3.63\pm 0.03}$
ViT-Base ViT-Small KDP (ours) + CODA-Prompt [30] $\mathbf{71.92\pm 0.50}$ $5.61\pm 0.34$ $\mathbf{83.72\pm 0.08}$ $6.40\pm 0.06$
$\varnothing$ ViT-Base + L2P [35] $73.94\pm 0.22$ $4.41\pm 0.18$ $83.02\pm 0.47$ $6.06\pm 0.47$
$\varnothing$ ViT-Base + DualPrompt [34] $73.18\pm 0.33$ $3.45\pm 0.32$ $84.66\pm 0.87$ $5.91\pm 0.34$
$\varnothing$ ViT-Base + CODA-Prompt [30] $76.42\pm 0.17$ $4.31\pm 0.18$ $86.16\pm 0.17$ $5.63\pm 0.25$
ViT-Large ViT-Base KDP (ours) + L2P [35] $76.91\pm 0.40$ $\mathbf{3.15\pm 0.39}$ $86.56\pm 0.22$ $4.97\pm 0.07$
ViT-Large ViT-Base KDP (ours) + DualPrompt [34] $76.06\pm 0.12$ $3.77\pm 0.38$ $86.92\pm 0.24$ $\mathbf{4.77\pm 0.58}$
ViT-Large ViT-Base KDP (ours) + CODA-Prompt [30] $\mathbf{78.62\pm 0.57}$ $3.46\pm 0.53$ $\mathbf{87.13\pm 0.09}$ $5.30\pm 0.06$
Table 2: Results of prompt distillation using KD token on the ImageNet-R dataset and the CIFAR-100 dataset with different teacher-student models and different continual learning models.
Table 1 summarizes the results of all the aforementioned knowledge distillation methods on the ImageNet-R and CIFAR100 datasets using CODA-Prompt as the continual learner. The complete results for the L2P and DualPrompt methods are provided in the supplementary material. When the teacher model is $\varnothing$, it indicates that there is no knowledge distillation. In the table, DeiT [31] refers to only using the KD token and the KD classifier without incorporating our KD prompts.
From the table, we can see that the logit distillation methods (KD [13], DKD [40]) and feature distillation methods (FitNets [28], ReviewKD [4]) fail to address the limitations discussed in Sec. 4.2. The distilled knowledge is fused into the CL prompt pool, making it difficult to transfer effectively across tasks. As observed in the results, although these four distillation methods can slightly improve the accuracy of the original student model, the improvements are not significant and the forgetting rate remains high. In many cases, the forgetting effect is even worse than that of the original student model. A high forgetting rate indicates that the overall accuracy improvement mainly comes from distillation benefits in the current task, while the influence of distillation on previous tasks is minimal. This further supports our analysis of the limitations in Sec. 4.2.
The DeiT method, using the distillation token and the KD classifier, separates the distilled knowledge from the CL prompt pool. This dual processing approach for CL and KD partially mitigates the interference between task learning and knowledge distillation using prompts. Our KDP method further enhances the separation and cross-task transfer of distilled knowledge through KD-Prompts, ultimately achieving the best overall performance on both datasets as shown in Table 1.
Table 2 presents the distillation results of our KDP method on the L2P, DualPrompt and CODA-Prompt models. This demonstrates the generalization ability of our KDP method in different Prompt-based CL approaches. All three prompt-based CL methods can significantly improve their performance with our KDP method. The combination of CODA-Prompt and our KDP method achieves state-of-the-art (SOTA) performance, particularly benefiting CL with ViT-Small as the backbone. It approaches the performance of larger models while significantly reducing inference computational costs.
5.5 Ablation Studies
The ablation study is divided into two main aspects. The first aspect explores the number of block layers in which prompts are inserted and the length of each prompt. The second aspect examines the impact of using the KD classifier and the corresponding KD token.
Multiple Layers & KD Prompt Length: For this study, we conducted experiments under the CODA-Prompt and L2P methods. The teacher model backbone is ViT-Base and the student model backbone is ViT-Small. We used the ImageNet-R dataset, where the number of tasks is set to 10. Note that a KD prompt in our model is denoted as $\mathbf{P}^{\text{kd}}\in\mathbb{R}^{L_{p}\times D}$. In the left-hand side of Fig. 3, we show the relationship between the length of the KD prompt, denoted as $L_{p}$, and the accuracy of the KDP model. In this case, the number of inserted layers of the ViT backbone is fixed at 12. In the right-hand side of Fig. 3, we present the relationship between the number of inserted block layers, denoted as $n$, and the accuracy. In this case, $L_{p}=6$.
From the results of the two plots, it can be observed that the length of KD prompts has no strong correlation with the final performance. Therefore, we select $L_{p}=6$, which performs better in CODA-Prompt. As the number of inserted layers $n$ increases, accuracy shows an upward trend in both methods, with a more pronounced effect in L2P. Based on this observation, we set the number of layers to $12$.
Figure 3: Left: The relationship between the length of KD prompt inserted into the model and accuracy. Right: The relationship between the number of layers where KD prompts are inserted and accuracy. Teacher: Vit-Base; Student: ViT-Small; Dataset: ImageNet-R.
KD Classifier: Our method by default adopts the KD classifier structure from DeiT [31] (including the corresponding KD token) to process the soft labels from the teacher. Here, we conducted an experiment to analyze the impact of using the KD classifier.
If neither KD prompts nor the KD classifier is added, the method corresponds to the normal KD [13] in logit distillation. If KD prompts are removed while only adding the KD classifier with the KD token, the method corresponds to DeiT [31]. If both components are incorporated, the method follows the default KDP setup as shown in Fig. 2. Additionally, we design an experiment where only KD prompts are added, in which case the student model has only a single classifier.
As shown in Table 3, the CODA-Prompt method is more heavily relying on the KD classifier for processing. The best performance is achieved when the KD prompts and the KD classifier are used together. However, for the L2P method, the KD classifier does not contribute significantly to improving the distillation. In fact, removing the KD classifier structure leads to an increase in accuracy. However, both CODA-Prompt and L2P depend on KD prompts to achieve better performance, which further validates the effectiveness of the KDP method.
Model KD-Prompts KD-Classifier Avg. Acc ($\uparrow$)
ViT-Small + CODA-Prompt [30] × × $69.91$
ViT-Small + CODA-Prompt [30] × $70.74$
ViT-Small + CODA-Prompt [30] × $70.21$
ViT-Small + CODA-Prompt [30] $\mathbf{71.92}$
ViT-Small + L2P [35] × × $63.97$
ViT-Small + L2P [35] × $64.99$
ViT-Small + L2P [35] × $\mathbf{69.00}$
ViT-Small + L2P [35] $68.18$
Table 3: Ablation study on Split ImageNet-R dataset. Comparison of different models with KD prompts and KD classifier.
6 Conclusion and Discussion
We introduce the problem of Continual Distillation Learning (CDL), which aims to improve prompt-based continual learning models using knowledge distillation. We first empirically studied logit distillation and feature distillation in the CDL setup, where three different prompt-based continual learning methods (L2P, DualPrompt and CODA-Prompt) are used. We found that these previous KD methods are not effective for the CDL problem. Therefore, we proposed a novel method, named Knowledge Distillation based on Prompts (KDP), to tackle the CDL problem. In KDP, a new type of prompt is introduced that is mainly designed for the distillation of knowledge. These additional learnable prompts significantly improve the learning performance of the student model in CDL. Experiments on two commonly used benchmarks for continual learning demonstrate the effectiveness of the proposed KDP method.
Limitations. CDL models require training a teacher model and a student model jointly. The total training time and memory consumption are increased.
Acknowledgments
This work was supported in part by the DARPA Perceptually-enabled Task Guidance (PTG) Program under contract number HR00112220005, the Sony Research Award Program, and the National Science Foundation (NSF) under Grant No. 2346528.
Supplementary Material
7 Why limiting the scope to prompt-based continual learning?
Model Accuracy Prompt-based?
iCaRL [26]-ResNet18 55.25 No
iCaRL [26]-ResNet34 56.65 No
LWF [19]-ResNet18 44.70 No
LWF [19]-ResNet34 41.46 No
L2P [35]-Tiny 60.68 Yes
L2P [35]-Small 77.71 Yes
L2P [35]-Base 83.02 Yes
L2P [35]-Large 86.36 Yes
CODA-Prompt [30]-Tiny 65.05 Yes
CODA-Prompt [30]-Small 82.18 Yes
CODA-Prompt [30]-Base 86.16 Yes
CODA-Prompt [30]-Large 88.97 Yes
Table 4: Comparison of accuracy on the CIFAR100 dataset between CL methods using different sizes of CNN backbones and CL methods using different sizes of ViT backbones. The number of tasks in continual learning is 10.
This is due to the following two reasons:
  • Prompt-based continual learning methods achieve state-of-the-art performance. In Table 4, we compare CODA-Prompt [30] and L2P [35] against iCaRL [26] and LWF [19] which are not prompt-based, we can clearly see that CODA-Prompt and L2P achieves better accuracy.
  • Traditional CNN-based continual learning models do not improve or have little improvement with larger backbones. In Table 4, we used ResNet18 and ResNet34 for iCaRL and LWF. The change of backbone results in 1.4% and -3.2% for iCaRL and LWF, respectively.
8 Logit Distillation Details
Logit distillation is one of the most classic forms of knowledge distillation. It aims to have a smaller student model learn the logits output of a larger or more accurate teacher model. In our CDL setup, we conducted experiments using two types of logit knowledge distillation methods: normal knowledge distillation (KD) [13] and Decoupled Knowledge Distillation (DKD) [40].
Normal Knowledge Distillation (KD): This approach was initially proposed by Hinton et al. [13], where distillation transfers the knowledge from a large, complex teacher model to a smaller student model, helping the latter to approximate the teacher model in terms of performance and accuracy. To achieve this, Hinton et al. designed a method where the logit output of the teacher model serves as “soft labels”, guiding the student model’s training. After passing through the softmax layer, the output values provide probabilities for each class.
$p_{i}=\frac{\exp(z_{i}/\tau)}{\sum_{j}\exp(z_{j}/\tau)},$ (11)
where $z_{i}$ is the logit, and $p_{i}$ is the predicted probability for class $i$. The soft target is the class probability distribution of teacher model. In the distillation process, a temperature parameter $\tau$, is introduced to smooth the output distribution of the model, making it easier for the student model to learn the subtle differences between classes.
In the teacher-student model shown in Fig. 2, the class tokens of the teacher model and the student model processed by the pre-trained ViT backbones are connected to the teacher classifier and student classifier to output their logits, respectively. The loss function of the student model for logit distillation is defined as:
$\mathcal{L}_{\text{S}}$ $=(1-\alpha)\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)+\alpha\mathcal{L}_{\text{KD}}+\lambda\mathcal{L}_{\text{pool}},$ (12)
$\mathcal{L}_{\text{KD}}$ $=\tau^{2}\sum_{i}p_{i}^{\mathcal{T}}\log\left(\frac{p_{i}^{\mathcal{T}}}{p_{i}^{\mathcal{S}}}\right),$ (13)
where $\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)$ represents the cross-entropy classification loss used for learning with true label $y$, $g_{\phi}$ is the student classifier, and $f_{b}$ denotes the pre-trained ViT backbone and we only use the final class token into the classifier. $\mathcal{L}_{\text{KD}}$ represents the knowledge distillation loss, i.e., the KL divergence between the teacher’s probability distribution $p_{i}^{\mathcal{T}}$ and the student’s probability distribution $p_{i}^{\mathcal{S}}$. The $\mathcal{L}_{\text{pool}}$ is a loss function specific to the prompt pool. Different prompt-based continual learning methods have their respective prompt pool loss functions. For example, please refer to the loss functions in Equations (2) and (4) for L2P and CODA-Prompt. $\alpha$ and $\lambda$ are hyperparameters used to balance the weights of loss components.
Decoupled Knowledge Distillation (DKD) [40]: Unlike traditional knowledge distillation, which uses a unified KL divergence to measure the difference between the outputs of the student model and the teacher model, DKD separates the logit distillation loss into target class and non-target class components. DKD considers that the target and non-target classes contain different information and should be handled separately during training. By decoupling these components, DKD allows the student model to better capture the confidence on the target class while learning the distribution of the non-target classes. The student model loss function is formulated as follows:
$\mathcal{L}_{\text{S}}$ $=(1-\alpha)\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)$
$+\alpha(\mathcal{L}_{\text{TCKD}}+\mathcal{L}_{\text{NCKD}})+\lambda\mathcal{L}_{\text{pool}},$ (14)
$\mathcal{L}_{\text{TCKD}}$ $=p_{t}^{\mathcal{T}}\log\left(\frac{p_{t}^{\mathcal{T}}}{p_{t}^{S}}\right)+p_{\setminus t}^{\mathcal{T}}\log\left(\frac{p_{t}^{\mathcal{T}}}{p_{\setminus t}^{\mathcal{S}}}\right)$ (15)
$\mathcal{L}_{\text{NCKD}}$ $=p_{\setminus t}^{\mathcal{T}}\sum_{\begin{subarray}{c}i\neq t\end{subarray}}\hat{p}_{i}^{\mathcal{T}}\log\left(\frac{\hat{p}_{i}^{\mathcal{T}}}{\hat{p}_{i}^{\mathcal{S}}}\right)$ (16)
where $\mathcal{L}_{\text{TCKD}}$ is the target class distillation loss, focused on aligning the student’s confidence with the teacher’s for the correct class. $\mathcal{L}_{\text{NCKD}}$ is the non-target class distillation loss, focused on matching the teacher and student model distributions for the incorrect classes. $[p_{t}^{\mathcal{T}},p_{\setminus t}^{\mathcal{T}}]$ represents the binary probabilities of the target class $p_{t}^{T}$ and all the other non-target classes $p_{\setminus t}^{\mathcal{T}}$ in teacher model, which can be calculated by Equation (11). $[p_{t}^{\mathcal{S}},p_{\setminus t}^{\mathcal{S}}]$ represents the binary probabilities in student model. Meanwhile, $\hat{p}_{i}^{\mathcal{T}}$ and $\hat{p}_{i}^{\mathcal{S}}$ are probability distributions among the non-target classes (without considering the $t$th class). Each element is calculated by:
$\hat{p}_{i}=\frac{\exp(z_{i}/\tau)}{\sum\limits_{j\neq t}\exp(z_{j}/\tau)}.$ (17)
Methods Split ImageNet-R Split CIFAR-100
Teacher Student Avg. Acc ($\uparrow$) Forgetting ($\downarrow$) Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
$\varnothing$ ViT-Small $63.82\pm 0.25$ $6.52\pm 0.31$ $77.71\pm 0.49$ $7.12\pm 0.33$
ViT-Base ViT-Small KD [13] $63.97\pm 0.62$ $6.51\pm 0.06$ $79.64\pm 0.04$ $6.35\pm 0.02$
ViT-Base ViT-Small DKD [40] $62.91\pm 0.27$ $6.55\pm 0.17$ $78.21\pm 0.12$ $9.13\pm 0.07$
ViT-Base ViT-Small FitNets [28] $64.29\pm 0.09$ $6.37\pm 0.17$ $79.56\pm 0.39$ $5.89\pm 0.36$
ViT-Base ViT-Small ReviewKD [4] $63.64\pm 0.34$ $6.36\pm 0.58$ $78.50\pm 0.39$ $8.04\pm 0.75$
ViT-Base ViT-Small DeiT [31] $64.99\pm 0.49$ $3.83\pm 0.85$ $79.56\pm 0.07$ $6.71\pm 0.16$
ViT-Base ViT-Small KDP (ours) $\mathbf{68.18\pm 0.03}$ $\mathbf{2.08\pm 0.28}$ $\mathbf{81.79\pm 0.66}$ $\mathbf{4.31\pm 0.27}$
$\varnothing$ ViT-Base $73.94\pm 0.22$ $4.41\pm 0.18$ $83.02\pm 0.47$ $6.06\pm 0.47$
ViT-Large ViT-Base KD [13] $74.12\pm 0.42$ $4.60\pm 0.55$ $85.00\pm 0.34$ $\mathbf{4.48\pm 0.51}$
ViT-Large ViT-Base DKD [40] $74.58\pm 0.01$ $4.69\pm 0.06$ $83.29\pm 0.24$ $4.99\pm 0.17$
ViT-Large ViT-Base FitNets [28] $70.39\pm 0.23$ $5.84\pm 0.06$ $83.60\pm 0.02$ $5.21\pm 0.71$
ViT-Large ViT-Base ReviewKD [4] $72.17\pm 0.26$ $6.11\pm 0.08$ $83.12\pm 0.65$ $7.97\pm 0.27$
ViT-Large ViT-Base DeiT [31] $73.99\pm 0.01$ $5.09\pm 0.02$ $84.21\pm 0.71$ $6.06\pm 0.93$
ViT-Large ViT-Base KDP (ours) $\mathbf{76.91\pm 0.40}$ $\mathbf{3.15\pm 0.39}$ $\mathbf{86.56\pm 0.22}$ $4.97\pm 0.07$
Table 5: The continual knowledge distillation results on the CIFAR-100 dataset and the ImageNet-R dataset with different teacher-student models based on L2P [35].
Methods Split ImageNet-R Split CIFAR-100
Teacher Student Avg. Acc ($\uparrow$) Forgetting ($\downarrow$) Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
$\varnothing$ ViT-Small $65.51\pm 0.11$ $5.93\pm 0.03$ $79.85\pm 0.57$ $6.12\pm 0.32$
ViT-Base ViT-Small KD [13] $65.68\pm 0.06$ $7.26\pm 0.29$ $80.16\pm 0.54$ $5.76\pm 0.18$
ViT-Base ViT-Small DKD [40] $65.44\pm 0.05$ $7.27\pm 0.23$ $80.44\pm 0.46$ $6.96\pm 0.44$
ViT-Base ViT-Small FitNets [28] $66.20\pm 0.14$ $5.93\pm 0.17$ $80.70\pm 0.17$ $5.73\pm 0.21$
ViT-Base ViT-Small ReviewKD [4] $65.69\pm 0.91$ $6.56\pm 0.46$ $80.33\pm 0.23$ $5.86\pm 0.53$
ViT-Base ViT-Small DeiT [31] $65.82\pm 0.48$ $4.00\pm 0.19$ $80.64\pm 0.32$ $5.67\pm 0.59$
ViT-Base ViT-Small KDP (ours) $\mathbf{68.77\pm 0.16}$ $\mathbf{3.13\pm 0.25}$ $\mathbf{81.78\pm 0.17}$ $\mathbf{3.63\pm 0.03}$
$\varnothing$ ViT-Base $73.18\pm 0.33$ $3.45\pm 0.32$ $84.66\pm 0.87$ $5.91\pm 0.34$
ViT-Large ViT-Base KD [13] $73.90\pm 0.14$ $\mathbf{3.31\pm 0.04}$ $84.67\pm 0.53$ $\mathbf{4.52\pm 0.55}$
ViT-Large ViT-Base DKD [40] $75.24\pm 0.33$ $4.15\pm 0.23$ $84.93\pm 0.16$ $4.95\pm 0.12$
ViT-Large ViT-Base FitNets [28] $71.23\pm 0.04$ $5.71\pm 0.55$ $83.12\pm 0.86$ $8.33\pm 1.36$
ViT-Large ViT-Base ReviewKD [4] $72.19\pm 0.01$ $5.72\pm 0.20$ $84.11\pm 0.92$ $5.19\pm 0.33$
ViT-Large ViT-Base DeiT [31] $76.03\pm 0.03$ $3.90\pm 0.01$ $85.73\pm 0.27$ $5.05\pm 0.43$
ViT-Large ViT-Base KDP (ours) $\mathbf{76.06\pm 0.12}$ $3.77\pm 0.38$ $\mathbf{86.92\pm 0.24}$ $4.77\pm 0.58$
Table 6: The continual knowledge distillation results on the CIFAR-100 dataset and the ImageNet-R dataset with different teacher-student models based on DualPrompt [34].
9 Feature Distillation Details
Feature distillation focuses on transferring the intermediate representations from a teacher model to a student model. It can leverage richer, layer-wise information within the teacher model to guide the student model’s learning. The student model can benefit from an understanding of feature relationships. In the CDL model in Fig. 2, the backbone ViT consists of multiple blocks, and we use the internal tokens outputted by each block as features. This allows the internal tokens of the student model to learn the information in the internal tokens of the teacher model. It is worth noting that during knowledge distillation, the backbone of the student model remains frozen while processing the internal tokens, which is a characteristic of prompt-based continual learning methods. In this paper, we build two methods based on the handling of internal tokens: FitNets [28] and Review Knowledge Distillation (ReviewKD) [4].
FitNets [28]: This method is one of the most classic feature distillation methods, aimed at adding a distillation loss to the intermediate layers. It uses the intermediate feature representations of the teacher model as hints to guide the student model’s learning. Here, we select the output of the last block of the teacher model as the hint. Similarly, the student model selects the output of the corresponding block for learning, constructing the feature distillation loss. Finally, the total loss function for the student model is
$\mathcal{L}_{\text{S}}$ $=\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)+\alpha\mathcal{L}_{\text{hint}}+\lambda\mathcal{L}_{\text{pool}},$ (18)
$\mathcal{L}_{\text{hint}}$ $=\left\|f_{b-1}^{\mathcal{T}}(\mathbf{x})-F_{M}(f_{b-1}^{\mathcal{S}}(\mathbf{x}))\right\|^{2},$ (19)
where $\mathcal{L}_{\text{hint}}$ is the feature distillation loss, calculated using the Mean Squared Error (MSE). $\alpha$ and $\lambda$ are used to balance the weights of loss components. $f_{b-1}$ indicates the feature output (internal tokens) after the last block of the ViT model. The student feature is transformed into the same size as the teacher feature with the mapping layer $F_{M}$, which is simply a fully-connected layer in the network.
ReviewKD [4]: It innovates by “reviewing” multiple hidden layers from both the teacher and student models, offering a more comprehensive approach to capturing hierarchical features across the entire model. It reviews the multiple layers utilizing the concept of residual learning. For instance, the feature from $n_{th}$ block of the student is aggregated with the feature from $(n-1)_{th}$ block of the student to mimic the feature from $(n-1)_{th}$ block of the teacher. The total loss function of the student in ReviewKD is
$\mathcal{L}_{\text{S}}=\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)+\alpha\mathcal{L}_{\text{RKD}}+\lambda\mathcal{L}_{\text{pool}},$ (20)
$\mathcal{L}_{\text{RKD}}=\mathcal{D}(\mathbf{F}^{\mathcal{S}}_{n},\mathbf{F}^{\mathcal{T}}_{n})+\sum_{j=n-1}^{1}\mathcal{D}\left(\mathcal{U}(\mathbf{F}_{j}^{\mathcal{S}},\mathbf{F}_{j+1}^{\mathcal{S}}),\mathbf{F}^{\mathcal{T}}_{j}\right),$ (21)
where $\mathcal{L}_{\text{RKD}}$ is the reviewKD loss, and $\mathcal{D}$ is L2 distance between the student features and teacher features. All student features in the equations have passed through the mapping layer $F_{M}$ to make them the same dimension as the teacher features. $\mathbf{F}_{j}^{\mathcal{S}}$ and $\mathbf{F}_{j}^{\mathcal{T}}$ are the features output by the student model and teacher model, respectively, after passing through $j$ blocks. $\mathbf{F}_{j+1}^{\mathcal{S}}$ represents the fused student features at the $(j+1)_{th}$ block. $\mathcal{U}$ is a module used to fuse features, which performs a weighted combination of the two input features. $n$ is the total number of blocks in the ViT backbone. Therefore, in the student model, the fused features obtained at each block are passed to the next higher block to form new feature fusion.
10 Continual Distillation Results
Tables 8 and 6 present the results of the aforementioned knowledge distillation methods on the ImageNet-R and CIFAR-100 datasets, using L2P [35] and DualPrompt [34] as the continual learners. The results indicate that the KD-Token method consistently optimizes and improves the performance of the original student model across different datasets and various continual learners. Notably, it achieves the highest accuracy in the L2P and DualPrompt models on the CIFAR-10 dataset. Additionally, on the ImageNet-R dataset, the KD-Token method attains the highest accuracy under the base-to-small scheme for both models.